What is the difference between a progressive versus regressive tax. Give examples of both?
Answer #1

In case of progressive tax, as taxable income increases (decreases), marginal and average tax rates increase (decrease). For example, consider the following tax structure:

When Taxable income is upto $10,000, Income tax rate = 5% and when Taxable income lies between $10,000 and $50,000, Income tax rate = 25%. Then,

For taxable income of $5,000, Tax paid = $5,000 x 5% = $250.

Marginal tax rate = $250/$5,000 = 5%

Average tax rate = $250/$5,000 = 5%

For taxable income of $40,000, Tax paid ($) = 10,000 x 5% + (50,000 - 40,000) x 25% = 500 + (10,000 x 25%)

= 500 + 2,500 = 3,000

Marginal tax rate = 25%

Average tax rate = $3,000/$40,000 = 0.075 = 7.5%.

Since, with increase in taxable income (from $5,000 to $40,000), marginal tax rate is higher (25% > 5%) and average tax rate is higher (7.5% > 5%), this is a progressive tax system.

In contrast, for regressive tax, as taxable income increases (decreases), marginal and average tax rates decrease (increase). For example, consider the following tax structure:

When Taxable income is upto $10,000, Income tax rate = 25% and when Taxable income lies between $10,000 and $50,000, Income tax rate = 5%. Then,

For taxable income of $5,000, Tax paid = $5,000 x 25% = $1,250.

Marginal tax rate = 25%

Average tax rate = $1,250/$5,000 = 25%

For taxable income of $40,000, Tax paid ($) = 10,000 x 25% + (50,000 - 40,000) x 55% = 2,500 + (10,000 x 5%)

= 2,500 + 500 = 3,000

Marginal tax rate = 5%

Average tax rate = $3,000/$40,000 = 0.075 = 7.5%.

Since, with increase in taxable income (from $5,000 to $40,000), marginal tax rate is lower (5% < 25%) and average tax rate is lower (25% < 5%), this is a regressive tax system.